OpenStudy (anonymous):

9x + 4= 7x - 8x + 16

OpenStudy (anonymous):

@iGreen

OpenStudy (zpupster):

combine like terms ex 7x-8x =?? then subtract 4 from both sides getting the numbers on one side do that then we can proceed

OpenStudy (anonymous):

7x = 20

OpenStudy (zpupster):

no 7x-8x =?? this equals -x subtracting 4 gives 9x = -x + 12

OpenStudy (zpupster):

now add x to both sides

OpenStudy (anonymous):

8x = 12

OpenStudy (zpupster):

no we are adding x to eliminate the x on the right hand side 9x = -x + 12 9x+x=(-x+x) + 12 10x= 12 now to get x by itself we divide both sides by 10 10x/10=12/10 the 10s cancel

OpenStudy (zpupster):

if you notice --if we do something to one side of the equation we must do it to the other